OpenStudy (anonymous):

Discrete math: Someone invests $3,000 at 12% annual interest, compounded quarterly. Let A[sub(n)] represent the amount after n years. Find recurrence relation for A[sub(0)], A[sub(1)], ...

OpenStudy (tkhunny):

Have you considered accumulating at 4% per quarter?

OpenStudy (anonymous):

so like An (n is sub) = 3000 * (1/4 * 0.12)n? how off am I on that but anyway wouldn't it be 3%/quarter since it's 12% annually 4 quarters (by def) in a yr so 12/4 = 3%... or am i even completely incapable of simple math...

OpenStudy (anonymous):

no it's gotta be n-1? no i don't know how to get a recursive relation itself from that i need an example.

OpenStudy (perl):

A(0) = 3000 , since we start at 3000

OpenStudy (perl):

The general formula for compounding interest n times a year is A(t) = P ( 1 + r/n) ^ (n*t)

OpenStudy (perl):

but you want a recurrence relation

OpenStudy (anonymous):

what's r? why 1 + ?

OpenStudy (perl):

r is the annual rate of interest , n is the number of times you compound a year

OpenStudy (perl):

this is a different n than what you are using

OpenStudy (anonymous):

oh gotcha..

OpenStudy (anonymous):

right yeah

OpenStudy (perl):

The general formula for compounding interest k times a year is A(t) = P ( 1 + r/k) ^ (k*t)

OpenStudy (perl):

:)

OpenStudy (anonymous):

ok thank you but what is P?

OpenStudy (perl):

P is the principal , here 3000

OpenStudy (anonymous):

So A(n) = P ( 1 + r/k ) ^(k*n) ) = 3000 (1 + 0.12/0.25) ^(0.25*n)?

OpenStudy (anonymous):

sure much appreciated

OpenStudy (perl):

oh wait, the directions say Let A[sub(n)] represent the amount after n years.

OpenStudy (perl):

so you are right, n is equal to t in the formula

OpenStudy (perl):

lets make a table

OpenStudy (anonymous):

yeah ok so then you don't need to take a different approach is what you're saying? lol my understanding is sufficient?

OpenStudy (anonymous):

let's

OpenStudy (perl):



OpenStudy (perl):